# DiCoDe: Diffusion-Compressed Deep Tokens for Autoregressive Video Generation with Language Models

### v1

@@ 1 Introduction
Figure 1: A VQ-VAE-based encoder [53] represent a 2-second $256\times 256$ video clip consisting of 16 frames with 16,384 low-level tokens. Harvesting the prior knowledge of video diffusion models, DiCoDe compresses the same video clip into 16 high-level deep tokens, achieving a 1000$\times$ reduction in token count. This extremely high compression rate facilitates chronological video modeling with autoregressive language models.
@@
In this work, we introduce DiCoDe, a novel approach that leverages Diffusion-Compressed Deep Tokens to generate videos autoregressively with a language model in a chronological and scalable manner. By leveraging the prior knowledge of the video diffusion model [49], DiCoDe learns a frame-level tokenizer to encode fixed-length video clips into high-level tokens, trained through clip-level denoising. In DiCoDe, a 2-second video clip is represented by 16 continuous tokens, which achieves a compression rate that is 1,000$\times$ higher than that of low-level discrete tokens. This significant reduction in token count makes it feasible to model videos temporally with AR language models. These learned deep tokens essentially serve as a “language” for videos and are designed to satisfy the following properties: 1) Temporally causal: By encoding video clips in a way that preserves temporal order, DiCoDe aligns with the sequential nature of AR models and video data; 2) Highly compressed: By leveraging the prior knowledge of video diffusion model, videos can be represented with a manageable number of tokens for efficient AR modeling; 3) Compatible with image data: Our frame-level tokenizer allows images to be effectively represented, alleviating the scarcity of high-quality video-text data.
@@ 3.1 Designing a Language for Videos
This “language”, i.e. deep tokens, serves as a proxy between the video diffusion model and the AR model. We justify the design of deep tokens from both cognitive and practical perspectives. The following principles must be adhered to: i) Sequential in time. To match the sequential nature of AR models and video data, deep tokens must be chronological, which is crucial for naturally extending to longer videos and enhancing scalability. ii) Highly compressed. While an adult reads 200-300 words per minute in English [4], a one-minute $256\times 256$ video of 8FPS typically requires 8FPS×256Tokens / $8\text{FPS}\times 256\text{Tokens / Frame}\times 60s\approx 120\text{k}$ tokens to model, which is obviously unmanageable for AR models. From a cognitive point of view, modeling videos analogy to language requires an extremely compressed tokenization method. iii) Compatible with image data. High-quality video-text data is scarce compared to image-text data, with commonly used datasets like WebVid [1] are significantly smaller in magnitude than image-text datasets such as LAION [33]. However, autoregressive models are data-hungry. To mitigate the scarcity of video-text data, deep tokens should integrate seamlessly with image data, allowing the model to leverage the abundant image-text data for much richer semantic information.
Given a video of varying length $\mathbf{V}\in\mathbb{R}^{T\times H\times W\times 3}$, DiCoDe first sample it with a fixed large frame rate $f$, achieving an $f\times$ reduction in temporal. The sampled video $\mathbf{V}\in\mathbb{R}^{T/f\times H\times W\times 3}$ is split into frames $\mathbf{F}_{0},\mathbf{F}_{f},\dots,\mathbf{F}_{T}$. The frames are then encoded into high-level semantic-rich tokens individually using an image encoder $\mathcal{E}$ as feature $\mathbf{Z}_{t}\in\mathbb{R}^{N\times C}$, where $t$ is the frame index, $N$ is the number of tokens, $C$ is the channel dimension. Despite $N$ is usually already smaller than the number of typical low-level tokens, it is still too large to fit into AR models if $T$ is large. The high-level tokens are further compressed into deep tokens using a query transformer $\mathcal{Q}$. Specifically, $\mathcal{Q}$ learn a fixed set of $N_{q}$ queries $\mathbf{q}\in\mathbb{R}^{N_{q}\times C}$ by concatenating $[\mathbf{q},\mathbf{Z}_{t}]$ and passing them through a multi-layer self-attention transformer. The final output of $\mathcal{Q}$ is a set of $N_{q}$ high-level continuous tokens $\mathbf{Q}_{t}\in\mathbb{R}^{N_{q}\times C}$, which is used as the deep tokens.
@@ 3.4 Explicitly Introducing the Variance
Due to the decoupled and offline essences of diffusion loss, the target of the AR model is deterministic. Unlike in discrete scenario where the target is a categorical distribution, the target in continuous scenario is a point estimate. As found in MAR [23], directly employing $\mathcal{L}_{2}$ loss will result in disastrous performance. An informal justification is that $\mathcal{L}_{2}$ loss implicitly assumes the target variable follows a Gaussian distribution with a fixed variance. Specifically, if the target is Gaussian and $p_{\text{high}}(q)$ $\sim\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ is the mean and $\sigma$ is the standard deviation, minimizing the $\mathcal{L}_{2}$ loss is equivalent to minimizing the negative log-likelihood of the target distributed, which is $\mathcal{L}_{\text{NLL}}=\frac{1}{2\sigma^{2}}||q-\mu||^{2}+\text{constant}$. Since no constraints are applied to the deep tokens, the assumption of a Gaussian distribution does not hold in our case.
@@ 5 Conclusion
Limitation.
The primary limitation comes from the data. DiCoDe excels with natural scenes and nonrigid objects such as ocean, mountain, and forest but sometimes struggles with rigid objects due to the WebVid-10M dataset’s focus on natural scenes. The length of WebVid-10M (17.5s in average) also heavily restricts the potential for long video generation. Another limitation is the design of deep tokens. While our assumption that video redundancy allows for compression into a few tokens holds in most cases, it may fail with videos that have sudden transitions.
@@ 6 Implementation Details
The autoregressive models are trained in a progressive manner for faster convergence. For image pre-training, it is first trained on image data listed in Tab. 1 for 100k iterations with a batch size of 256 using $\mathcal{L}_{2}$ Loss, then another 100k iterations using GMM Loss. For video training, it is first trained on mixed image and video data for 100k iterations with a batch size of 256 and image:video ratio of 1:1. Then it is trained on filtered motion videos for another 20k iterations. To filter the motion videos, we compute the optical flow between consecutive frames and 1 fps and average the flow magnitude as the motion score. We choose videos with motion score between 0.5 and 1.38, length between 556 and 4000 frames, resulting in 1M videos in total. The length of the video is limited to 256 frames, with a max context length of $80+(1+16)*16=352$, in which the [BOV] token is also counted. The videos are also filtered with a minimum aspect ratio of 0.3333, minimum resolution 200. To adapt the video diffusion models to the variance introduced by the autoregressive language models, further fine-tuning the video diffusion models for another 20k iterations based on the prediction of the autoregressive language models. For all training, we use a learning rate of 1e-4, cosine schedule with warm-up for 1k iterations, AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=1e-6$, weight decay $0.05$.

### v2

@@ Abstract
This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000.
@@ 1 Introduction
[…]
@@
[…]
@@ 3.1 Designing a Language for Videos
This “language”, i.e. deep tokens, serves as a proxy between the video diffusion model and the AR model. We justify the design of deep tokens from both cognitive and practical perspectives. The following principles must be adhered to: i) Sequential in time. To match the sequential nature of AR models and video data, deep tokens must be chronological, which is crucial for naturally extending to longer videos and enhancing scalability. ii) Highly compressed. While an adult reads 200-300 words per minute in English [4], a one-minute $256\times 256$ video of 8FPS typically requires $8\text{FPS}\times 256\text{Tokens / Frame}\times 60s\approx 120\text{k}$ tokens to model, which is obviously unmanageable for AR models. From a cognitive point of view, modeling videos analogy to language requires an extremely compressed tokenization method. iii) Compatible with image data. High-quality video-text data is scarce compared to image-text data, with commonly used datasets like WebVid [1] are significantly smaller in magnitude than image-text datasets such as LAION [33]. However, autoregressive models are data-hungry. To mitigate the scarcity of video-text data, deep tokens should integrate seamlessly with image data, allowing the model to leverage the abundant image-text data for much richer semantic information.
Given a video of varying length $\mathbf{V}\in\mathbb{R}^{T\times H\times W\times 3}$, DiCoDe first sample it with a fixed large frame rate $f$, achieving an $f\times$ reduction in temporal. The sampled video $\mathbf{V}\in\mathbb{R}^{T/f\times H\times W\times 3}$ is split into frames $\mathbf{F}_{0},\mathbf{F}_{f},\dots,\mathbf{F}_{T}$. The frames are then encoded into high-level semantic-rich tokens individually using an image encoder $\mathcal{E}$ as feature $\mathbf{Z}_{t}\in\mathbb{R}^{N\times C}$, where $t$ is the frame index, $N$ is the number of tokens, $C$ is the channel dimension. Despite $N$ is usually already smaller than the number of typical low-level tokens, it is still too large to fit into AR models if $T$ is large. The high-level tokens are further compressed into deep tokens using a query transformer $\mathcal{Q}$. Specifically, $\mathcal{Q}$ learn a fixed set of $N_{q}$ queries $\mathbf{q}\in\mathbb{R}^{N_{q}\times C}$ by concatenating $\mathbf{q},\mathbf{Z}_{t}]$ and passing them through a multi-layer self-attention transformer. The final output of $\mathcal{Q}$ is a set of $N_{q}$ high-level continuous tokens $\mathbf{Q}_{t}\in\mathbb{R}^{N_{q}\times C}$, which is used as the deep tokens.
@@ 3.4 Explicitly Introducing the Variance
[…]
@@ 5 Conclusion
Limitation. The primary limitation comes from the data. DiCoDe excels with natural scenes and nonrigid objects such as ocean, mountain, and forest but sometimes struggles with rigid objects due to the WebVid-10M dataset’s focus on natural scenes. The length of WebVid-10M (17.5s in average) also heavily restricts the potential for long video generation. Another limitation is the design of deep tokens. While our assumption that video redundancy allows for compression into a few tokens holds in most cases, it may fail with videos that have sudden transitions.
@@ 6 Implementation Details
The autoregressive models are trained in a progressive manner for faster convergence. For image pre-training, it is first trained on image data listed in Tab. 1 for 100k iterations with a batch size of 256 using $\mathcal{L}_{2}$ Loss, then another 100k iterations using GMM Loss. For video training, it is first trained on mixed image and video data for 100k iterations with a batch size of 256 and image:video ratio of 1:1. Then it is trained on filtered motion videos for another 20k iterations. To filter the motion videos, we compute the optical flow between consecutive frames and 1 fps and average the flow magnitude as the motion score. We choose videos with motion score between 0.5 and 1.38, length between 556 and 4000 frames, resulting in 1M videos in total. The length of the video is limited to 256 frames, with a max context length of 80+(1+16)16=35280+(1+16)*16=352, in which the [BOV] token is also counted. The videos are also filtered with a minimum aspect ratio of 0.3333, minimum resolution 200. To adapt the video diffusion models to the variance introduced by the autoregressive language models, further fine-tuning the video diffusion models for another 20k iterations based on the prediction of the autoregressive language models. For all training, we use a learning rate of 1e-4, cosine schedule with warm-up for 1k iterations, AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=1e-6$, weight decay $0.05$.